Calculate the equilibrium constant, Kc
for the following reaction
if at 350 K the equilibrium mixture contains 2.50 mol of N2, 3.50 mol of O2 and 0.06 mol of NO2 in a 2.0 dm3 vessel. (C3)

N2(g) + 2O2(g) = 2NO2(g)
1
Expert's answer
2020-11-20T10:03:40-0500

[N2] = 1.25 M

[O2] = 1.75 M

[NO2] = 0.03 M

"K_c = \frac{[NO_2]^2}{[N_2][O_2]}"

"K_c = \frac{(0.03)^2}{1.25 \times 1.75} = 4.1 \times 10^{-4}"